# Large Language Model Supply Chain: A Research Agenda

### v2

@@ 5.2. Model: Testing & Evaluation
Current State. The current state of LLM testing and evaluation, while rapidly evolving, still falls short of a holistic approach, with practices often fragmented and heavily reliant on standardized benchmarks that may not fully capture real-world performance or ethical considerations (Sun et al., 2024c; Chao et al., 2024; Chang et al., 2024; Liu et al., 2024c; Cao et al., 2024). This landscape can be broadly categorized into three critical dimensions: helpfulness, honesty, and harmlessness, each presenting unique challenges. Helpfulness testing primarily relies on benchmark datasets for tasks such as question answering, task completion, and knowledge retrieval (Sun et al., 2024a; Guo et al., 2023a; Kandpal et al., 2023), but often falls short in capturing real-world complexities, particularly in domains like code generation (Liu et al., 2024c; Du et al., 2024). The issue of data contamination further complicates performance assessment (Deng et al., 2023; Magar and Schwartz, 2022). Honesty testing focuses on factual correctness and consistency, employing benchmarks like TruthfulQA (Lin et al., 2021) and FELM (Zhao et al., 2024d), yet struggles with detecting subtle misinformation and balancing hallucination mitigation with generative capabilities (Ji et al., 2023a; Zhang et al., 2023b). Harmlessness testing encompasses a wide range of evaluations for biases, security vulnerabilities, and potential misuse (Gehman et al., 2020; Huang et al., 2023; Zhao et al., 2024b), but faces challenges in capturing context-dependent ethical considerations and anticipating novel misuse scenarios (Bhatt et al., 2023; Zhang et al., 2023a).
@@
Challenge I [SE Perspective]: Functionality and Performance Evaluation. Evaluating the helpfulness of LLMs is a critical aspect of ensuring their practical utility and widespread adoption. To this end, researchers have been developing benchmark datasets and tasks that measure LLM performance on capabilities such as question answering (Sun et al., 2024a; Bian et al., 2021; Zhao et al., 2024d), task completion (Guo et al., 2023a; Valmeekam et al., 2024; Efrat et al., 2022; Dong et al., 2023; Li et al., 2023b), and knowledge retrieval (Kandpal et al., 2023; Shi et al., 2024b; Gu et al., 2024) across diverse domains. These benchmarks not only test for general knowledge (Kandpal et al., 2023; Shi et al., 2024b; Gu et al., 2024) but also probe domain-specific expertise (Sun et al., 2024a; Guo et al., 2023a; Valmeekam et al., 2024), allowing for a comprehensive assessment of an LLM’s ability to provide useful and relevant outputs. However, there are still several formidable challenges that highlight not only the complexity inherent in measuring the utility of such models but also underscore the necessity for ongoing refinement in our approaches to evaluation.
@@
  • Opportunity: Developing Comprehensive Metrics & Benchmarks. First and foremost, the opportunity to develop more comprehensive metrics and benchmarks provides a pathway to better understand the performance of LLMs (Liu et al., 2024c; Yu et al., 2024a; McIntosh et al., 2024). Traditional benchmarks, while useful, often fail to capture the multifaceted nature of tasks LLMs are expected to perform, especially in areas like code generation (Liu et al., 2024c; Du et al., 2024; Yu et al., 2024a). While some repository-level datasets have been proposed to address this issue (Wang et al., 2024a; Liu et al., 2024d), they remain limited in scope and do not fully encompass the complexities involved in generating code at the project level. This highlights a significant gap in current benchmarking efforts, indicating a need for more robust metrics that can evaluate an LLM’s ability to understand project-specific contexts, manage dependencies across multiple files, and ensure consistency within a larger codebase. Developing such metrics remains a critical area for research and innovation.
@@
  • Opportunity: Mitigating Data Contamination. The issue of data contamination (Deng et al., 2023; Magar and Schwartz, 2022; Cao et al., 2024) significantly complicates the evaluation of LLMs. Data contamination occurs when a model is inadvertently exposed to information from the test set during training, leading to inflated performance metrics that do not accurately represent the model’s true capabilities (Li and Flanigan, 2024; Ponnuru et al., 2024). This challenge is particularly acute in domains like code generation (Dong et al., 2024a; Cao et al., 2024), where the vast amount of publicly available code means that models might “learn” specific solutions during training that they later reproduce during testing. Such instances of data contamination not only overestimate the model’s performance but also obscure our understanding of its ability to generate innovative solutions to new problems. While there have been efforts to quantify and detect data contamination (Li et al., 2024b; Golchin and Surdeanu, 2023b; Li, 2023; Golchin and Surdeanu, 2023a), effectively addressing this issue remains a challenge (Dekoninck et al., 2024; Balloccu et al., 2024; Oren et al., 2023; Cao et al., 2024). Opportunities in identifying and mitigating the impact of data contamination include the development of novel evaluation frameworks that can detect when a model is reproducing rather than generating solutions (Golchin and Surdeanu, 2023b; Li, 2023; Golchin and Surdeanu, 2023a). Moreover, there is a demand for advanced perturbation techniques (Kargupta et al., 2003) that preserve the integrity of the code while adding enough variability to prevent contamination. This involves developing flexible perturbation frameworks specifically designed for semantic class code evaluation datasets, capable of incorporating different approaches such as code refactoring and transformations that maintain logical equivalence. Additionally, developing testing metrics and benchmarks specifically designed to prevent data contamination will be crucial in ensuring reliable evaluations of LLM performance (Dong et al., 2024a; Jain et al., 2024; Li et al., 2024b).
@@
Challenge II [SE Perspective]: Reliability and Truthfulness Assessment. As LLMs become increasingly influential in various domains, ensuring their honesty and truthfulness is paramount to building trust and preventing the spread of misinformation. Honesty testing (Lin et al., 2021; Zhao et al., 2024d; Liang et al., 2022; Li et al., 2023a; Lee et al., 2022; Muhlgay et al., 2023) for LLMs involves assessing whether the models can consistently provide information that is not only factually correct but also free from deception or misleading implications. These tests aim to identify instances of hallucinated (Ji et al., 2023a; Zhang et al., 2023b; Agarwal et al., 2024) or fabricated information (Lin et al., 2021; Lee et al., 2022; Muhlgay et al., 2023) in LLM outputs, which can undermine their trustworthiness. Assessing the consistency and coherence of LLM outputs across multiple queries and prompts can reveal potential inconsistencies or contradictions, which may indicate a lack of factual grounding or honesty.
@@
  • Opportunity: Hallucination Mitigation. Hallucination mitigation in LLMs is an area of significant concern, with various innovative techniques employed to address the issue (Tonmoy et al., 2024; Gunjal et al., 2024; Alshahwan et al., 2024). These methods range from retrieval augmented generation (Gao et al., 2022; Peng et al., 2023; Varshney et al., 2023) to self-refinement through feedback and reasoning (Si et al., 2022; Ji et al., 2023b), each targeting different aspects of hallucination to ensure the accuracy and reliability of LLM outputs. However, there’s an open problem in balancing mitigation efforts with the preservation of LLMs’ generative capabilities, avoiding over-restriction that could stifle their performance. The development of LLMs with inherent mechanisms to prevent hallucinations is an exciting avenue, potentially leading to inherently more honest models.
@@
Challenge III [S&P + SE Perspective]: Safety and Ethical Compliance Evaluation. The challenge of evaluating LLMs for potential harmful outputs and behaviors is multifaceted, rooted in the need to detect and mitigate a broad spectrum of potential harms. Researchers have been developing benchmarks (Gehman et al., 2020; Cui et al., 2024; Huang et al., 2023; Ji et al., 2024; Nagireddy et al., 2024) that probe LLMs for the presence of harmful biases, stereotypes, or discriminatory language across various sensitive topics and demographic groups. Furthermore, testing LLMs for potential vulnerabilities to adversarial attacks (Zhao et al., 2024b; Wang et al., 2023; Qi et al., 2024),jailbreaks (Zhou et al., 2024; Deng et al., 2024; Chu et al., 2024a; Liu et al., 2023c; Dong et al., 2024b), or misuse (Bhatt et al., 2023; Zhang et al., 2023a; Fu et al., 2023; Hazell, 2023; Yao et al., 2024; Madani, 2023) by attackers ensures their outputs do not enable harmful actions or security breaches. Yet they are beset with the inherent challenge of predicting and counteracting the myriad ways in which these sophisticated models might be exploited or go awry.
@@
  • Opportunity: Advanced Security Testing and Mitigation. The domain of harmlessness testing for LLMs presents substantial opportunities to enhance the security and robustness of these systems. Developing sophisticated jailbreak prevention techniques (Xu et al., 2022; Zeng et al., 2024; Yi et al., 2024) can significantly improve LLM resilience against malicious attempts to bypass safety constraints. Furthermore, the implementation of comprehensive LLM red teaming protocols (Ge et al., 2023; Xu et al., 2024; Mazeika et al., 2024) offers a proactive approach to identifying and mitigating potential vulnerabilities before deployment. Such efforts can involve simulating various attack scenarios, from subtle prompt injections to more complex multi-turn manipulation attempts, thereby uncovering hidden weaknesses in LLM systems. Additionally, advancing adversarial testing methodologies (Zhao et al., 2024b; Wang et al., 2023; Zou et al., 2023b) can lead to the development of more robust models capable of maintaining safe and ethical outputs even under targeted attacks.
@@ 5.3. Model: Release & Sharing
Current State. The current state of model release and sharing in the LLM supply chain, while evolving rapidly, still falls short of the envisioned ideal. Platforms like Hugging Face have significantly improved accessibility and collaboration in model sharing. However, the management of supply chain risks remains a critical challenge, particularly in terms of model provenance and security assurances (Kathikar et al., 2023). Model cards and associated documentation, while intended to provide comprehensive information, often fail to accurately reflect the true nature and capabilities of models (Ajibode et al., 2024; Pepe et al., 2024; Liang et al., 2024; Jiang et al., 2023). This weak model provenance leaves the ecosystem vulnerable to malicious model poisoning and other forms of tampering (Zhao et al., 2024c; Daniel Huynh, Jade Hardouin, 2023). The reuse and adaptation of models through techniques like fine-tuning introduce complex model dependencies that can lead to risk propagation. However, the open-source model ecosystem currently lacks robust frameworks for modeling and governing these interdependencies. Models are essentially treated as black boxes, and unlike traditional open-source software, static inspection offers limited security assurances. Vulnerable pre-trained models may harbor hidden biases, backdoors, or other malicious features that evade standard safety evaluations (Abdali et al., 2024; Zhao et al., 2024c; Daniel Huynh, Jade Hardouin, 2023). License management for LLMs remains a contentious issue, as exemplified by the ongoing debates surrounding the licensing terms of models like Llama (Charlie Hull, 2023; Sriram Parthasarathy, 2024). The field is grappling with balancing open collaboration with intellectual property concerns and responsible AI development. Furthermore, the security of model hubs themselves has come under scrutiny. Services like model conversion tools hosted on platforms such as Hugging Face have been shown to be vulnerable to manipulation (Eoin Wickens, Kasimir Schulz, 2024), potentially allowing the introduction of malicious code into LLMs. The popularity of collaborative model merging techniques, while fostering innovation, also presents opportunities for introducing vulnerabilities (Hammoud et al., 2024).
@@
Challenge II [S&P + SE Perspective]: Risk Propagation in LLM Supply Chains. A significant challenge in the LLM supply chain lies in quantifying the propagation of risks across different models and components. The complex nature of LLMs, with their intricate architectures and numerous dependencies, makes it difficult to measure how vulnerabilities such as privacy breaches, biases, hallucinations, or potential backdoors propagate from one model to another. This challenge is exacerbated by the lack of standardized metrics and methodologies for assessing risk transmission. Model reuse techniques (Jiang et al., 2023), including fine-tuning, transfer learning, and distillation, add further complexity to this issue. When models are reused, there is a risk that vulnerabilities inherent in the original model can be unintentionally transferred to new iterations or applications. For instance, if a base model contains biases, these biases may be amplified or altered during the distillation process, leading to unforeseen consequences in downstream applications. Quantifying how such biases propagate through different stages of model reuse remains an open problem. Similarly, measuring the extent to which privacy vulnerabilities in training data persist through model iterations and affect derived models is not straightforward. The challenge is not only to identify these vulnerabilities but also to understand how they evolve as models are reused.
  • Opportunity: Developing Model Bill of Materials (MBOM). There lies a significant opportunity to enhance the security and integrity of LLMs through the development of standardized practices for generating and maintaining a Model Bill of Materials (MBOM) for pre-trained models (David Mor-Ofek, 2024), mirroring the concept of SBOM. Such standardization would improve supply chain transparency, enabling stakeholders to more effectively identify, assess, and mitigate risks. Moreover, fostering collaboration among researchers, industry practitioners, and regulatory bodies can lead to the establishment of robust best practices and guidelines for the responsible release and sharing of models. This collaborative approach would not only enhance the trustworthiness and accountability of LLMs across the supply chain but also ensure that risk mitigation strategies are holistic, timely, and aligned with evolving ethical and security standards, ultimately leading to a safer and more reliable LLM ecosystem.
@@
Challenge IV [S&P Perspective]: Pre-trained Model Code Poisoning/Backdoor. The proliferation of model hubs as centralized platforms for sharing and distributing pre-trained models introduces significant security challenges. These hubs, while fostering collaboration and accessibility, also become potential points of failure and attractive targets for malicious actors. Key challenges include ensuring the integrity and authenticity of uploaded models, preventing the distribution of models with embedded malware or backdoors (Zhao et al., 2024c), and protecting against unauthorized access or data breaches. The dynamic nature of these platforms, with frequent updates and contributions from various sources, further complicates maintaining consistent security standards. Moreover, the potential for supply chain attacks, where compromised models could affect numerous downstream applications and users, presents a systemic risk to the LLM ecosystem.
  • Opportunity: Advanced Security Solutions for Model Hub Integrity. The security challenges posed by model hubs open up significant opportunities for developing robust security frameworks and technologies. There is potential to create advanced model verification systems that can automatically scan for malware, backdoors, or other vulnerabilities in uploaded models (Zhao et al., 2024c). Opportunities exist to implement measures to reduce model poisoning attack vectors. This could include developing sophisticated typo-squatting detection mechanisms to prevent malicious actors from exploiting similar or deregistered model names, organizations, or namespaces, addressing both typo-squatting and AIJacking vulnerabilities (Adrian Wood, 2023; Nadav Noy, 2023). Enhanced token management systems could mitigate the risk of API token leaks from prominent organizations (Bar Lanyado, 2023), thereby preventing unauthorized access and potential supply chain attacks. These advancements would not only enhance the security of individual model hubs but also contribute to building a more resilient and trustworthy AI development ecosystem, reducing the risk of widespread supply chain attacks and unauthorized model access or manipulation (Eoin Wickens, Kasimir Schulz, 2024).

### v3

@@ 5.2. Model: Testing & Evaluation
Current State. The current state of LLM testing and evaluation, while rapidly evolving, still falls short of a holistic approach, with practices often fragmented and heavily reliant on standardized benchmarks that may not fully capture real-world performance or ethical considerations (Sun et al., 2024c; Chao et al., 2024; Chang et al., 2024; Liu et al., 2023f; Cao et al., 2024). This landscape can be broadly categorized into three critical dimensions: helpfulness, honesty, and harmlessness, each presenting unique challenges. Functionality and performance evaluation (helpfulness) primarily relies on benchmark datasets for tasks such as question answering, task completion, and knowledge retrieval (Sun et al., 2024a; Guo et al., 2023a; Kandpal et al., 2023), but often falls short in capturing real-world complexities, particularly in domains like code generation (Liu et al., 2023f; Du et al., 2024). The issue of data contamination further complicates performance assessment (Deng et al., 2024b; Magar and Schwartz, 2022). Reliability and truthfulness assessment (honesty) focuses on factual correctness and consistency, employing benchmarks like TruthfulQA (Lin et al., 2022) and FELM (Chen et al., 2023), yet struggles with detecting subtle misinformation and balancing hallucination mitigation with generative capabilities (Ji et al., 2023a; Zhang et al., 2023b). Safety and ethical compliance evaluation (harmlessness) encompasses a wide range of evaluations for biases, security vulnerabilities, and potential misuse (Gehman et al., 2020; Huang et al., 2023; Zhao et al., 2023), but faces challenges in capturing context-dependent ethical considerations and anticipating novel misuse scenarios (Bhatt et al., 2023; Zhang et al., 2023a).
@@
Challenge I [SE Perspective]: Functionality and Performance Evaluation. Evaluating the helpfulness of LLMs is a critical aspect of ensuring their practical utility and widespread adoption. To this end, researchers have been developing benchmark datasets and tasks that measure LLM performance on capabilities such as question answering (Sun et al., 2024a; Bian et al., 2021; Chen et al., 2023), task completion (Guo et al., 2023a; Valmeekam et al., 2023; Efrat et al., 2023; Dong et al., 2023; Li et al., 2023b), and knowledge retrieval (Kandpal et al., 2023; Shi et al., 2024c; Gu et al., 2024) across diverse domains. These benchmarks not only test for general knowledge (Kandpal et al., 2023; Shi et al., 2024c; Gu et al., 2024) but also probe domain-specific expertise (Sun et al., 2024a; Guo et al., 2023a; Valmeekam et al., 2023), allowing for a comprehensive assessment of an LLM’s ability to provide useful and relevant outputs. However, there are still several formidable challenges that highlight not only the complexity inherent in measuring the utility of such models but also underscore the necessity for ongoing refinement in our approaches to evaluation.
@@
  • Opportunity: Developing Comprehensive Metrics & Benchmarks. First and foremost, the opportunity to develop more comprehensive metrics and benchmarks provides a pathway to better understand the performance of LLMs (Liu et al., 2023f; Yu et al., 2024; McIntosh et al., 2024). Traditional benchmarks, while useful, often fail to capture the multifaceted nature of tasks LLMs are expected to perform, especially in areas like code generation (Liu et al., 2023f; Du et al., 2024; Yu et al., 2024). While some repository-level datasets have been proposed to address this issue (Wang et al., 2024a; Liu et al., 2024d), they remain limited in scope and do not fully encompass the complexities involved in generating code at the project level. This highlights a significant gap in current benchmarking efforts, indicating a need for more robust metrics that can evaluate an LLM’s ability to understand project-specific contexts, manage dependencies across multiple files, and ensure consistency within a larger codebase. Developing such metrics remains a critical area for research and innovation.
@@
  • Opportunity: Mitigating Data Contamination. The issue of data contamination (Deng et al., 2024b; Magar and Schwartz, 2022; Cao et al., 2024) significantly complicates the evaluation of LLMs. Data contamination occurs when a model is inadvertently exposed to information from the test set during training, leading to inflated performance metrics that do not accurately represent the model’s true capabilities (Li and Flanigan, 2024; Ponnuru et al., 2024). This challenge is particularly acute in domains like code generation (Dong et al., 2024a; Cao et al., 2024), where the vast amount of publicly available code means that models might “learn” specific solutions during training that they later reproduce during testing. Such instances of data contamination not only overestimate the model’s performance but also obscure our understanding of its ability to generate innovative solutions to new problems. While there have been efforts to quantify and detect data contamination (Li et al., 2024b; Golchin and Surdeanu, 2024; Li, 2023; Golchin and Surdeanu, 2023), effectively addressing this issue remains a challenge (Dekoninck et al., 2024; Balloccu et al., 2024; Oren et al., 2024; Cao et al., 2024). Opportunities in identifying and mitigating the impact of data contamination include the development of novel evaluation frameworks that can detect when a model is reproducing rather than generating solutions (Golchin and Surdeanu, 2024; Li, 2023; Golchin and Surdeanu, 2023). Moreover, there is a demand for advanced perturbation techniques (Kargupta et al., 2003) that preserve the integrity of the code while adding enough variability to prevent contamination. This involves developing flexible perturbation frameworks specifically designed for semantic class code evaluation datasets, capable of incorporating different approaches such as code refactoring and transformations that maintain logical equivalence. Additionally, developing testing metrics and benchmarks specifically designed to prevent data contamination will be crucial in ensuring reliable evaluations of LLM performance (Dong et al., 2024a; Jain et al., 2024a; Li et al., 2024b).
@@
Challenge II [SE Perspective]: Reliability and Truthfulness Assessment. As LLMs become increasingly influential in various domains, ensuring their honesty and truthfulness is paramount to building trust and preventing the spread of misinformation. Reliability and truthfulness assessment (Lin et al., 2022; Chen et al., 2023; Liang et al., 2023; Li et al., 2023a; Lee et al., 2022; Muhlgay et al., 2024) for LLMs involves assessing whether the models can consistently provide information that is not only factually correct but also free from deception or misleading implications. These tests aim to identify instances of hallucinated (Ji et al., 2023a; Zhang et al., 2023b; Agarwal et al., 2024) or fabricated information (Lin et al., 2022; Lee et al., 2022; Muhlgay et al., 2024) in LLM outputs, which can undermine their trustworthiness. Assessing the consistency and coherence of LLM outputs across multiple queries and prompts can reveal potential inconsistencies or contradictions, which may indicate a lack of factual grounding or honesty.
@@
  • Opportunity: Hallucination Mitigation. Hallucination mitigation in LLMs is an area of significant concern, with various innovative techniques employed to address the issue (Tonmoy et al., 2024; Gunjal et al., 2024; Alshahwan et al., 2024). These methods range from retrieval augmented generation (Gao et al., 2023; Peng et al., 2023; Varshney et al., 2023) to self-refinement through feedback and reasoning (Si et al., 2023; Ji et al., 2023d), each targeting different aspects of hallucination to ensure the accuracy and reliability of LLM outputs. However, there’s an open problem in balancing mitigation efforts with the preservation of LLMs’ generative capabilities, avoiding over-restriction that could stifle their performance. The development of LLMs with inherent mechanisms to prevent hallucinations is an exciting avenue, potentially leading to inherently more honest models.
@@
Challenge III [S&P + SE Perspective]: Safety and Ethical Compliance Evaluation. The challenge of evaluating LLMs for harmful outputs and behaviors is multifaceted. Researchers have been developing benchmarks (Gehman et al., 2020; Cui et al., 2024; Huang et al., 2023; Ji et al., 2023b; Nagireddy et al., 2024) that probe LLMs for the presence of harmful biases, stereotypes, or discriminatory language across various sensitive topics and demographic groups. Furthermore, testing LLMs for potential vulnerabilities to adversarial attacks (Zhao et al., 2023; Wang et al., 2023; Qi et al., 2024),jailbreaks (Zhou et al., 2024; Deng et al., 2024a; Chu et al., 2024a; Liu et al., 2023c; Dong et al., 2024b), or misuse (Bhatt et al., 2023; Zhang et al., 2023a; Fu et al., 2023; Hazell, 2023; Yao et al., 2023; Madani, 2023) by attackers ensures their outputs do not enable harmful actions or security breaches. Yet they are beset with the inherent challenge of predicting and counteracting the myriad ways in which these sophisticated models might be exploited or go awry.
@@
  • Opportunity: Security Testing. The domain of harmlessness testing for LLMs presents substantial opportunities to enhance the security and robustness of these systems. Developing sophisticated jailbreak prevention techniques (Xu et al., 2022; Zeng et al., 2024; Yi et al., 2024) can significantly improve LLM resilience against malicious attempts to bypass safety constraints. Furthermore, the implementation of comprehensive LLM red teaming protocols (Ge et al., 2024; Xu et al., 2024; Mazeika et al., 2024) offers a proactive approach to identifying and mitigating potential vulnerabilities before deployment. Such efforts can involve simulating various attack scenarios, from subtle prompt injections to more complex multi-turn manipulation attempts, thereby uncovering hidden weaknesses in LLM systems. Additionally, advancing adversarial testing methodologies (Zhao et al., 2023; Wang et al., 2023; Zou et al., 2023b) can lead to the development of more robust models capable of maintaining safe and ethical outputs even under targeted attacks.
@@ 5.3. Model: Release & Sharing
Current State. The current state of model release and sharing in the LLM supply chain, while evolving rapidly, still falls short of the envisioned ideal. Platforms like Hugging Face have significantly improved accessibility and collaboration in model sharing. However, the management of supply chain risks remains a critical challenge, particularly in terms of model provenance and security assurances (Kathikar et al., 2023). Model cards and associated documentation, while intended to provide comprehensive information, often fail to accurately reflect the true nature and capabilities of models (Ajibode et al., 2024; Pepe et al., 2024; Liang et al., 2024; Jiang et al., 2023). This weak model provenance leaves the ecosystem vulnerable to malicious model poisoning and other forms of tampering (Zhao et al., 2024b; Daniel Huynh, Jade Hardouin, 2023). The reuse and adaptation of models through techniques like fine-tuning introduce complex model dependencies that can lead to risk propagation. However, the open-source model ecosystem currently lacks robust frameworks for modeling and governing these interdependencies. Models are essentially treated as black boxes, and unlike traditional open-source software, static inspection offers limited security assurances. Vulnerable pre-trained models may harbor hidden biases, backdoors, or other malicious features that evade standard safety evaluations (Abdali et al., 2024; Zhao et al., 2024b; Daniel Huynh, Jade Hardouin, 2023). License management for LLMs remains a contentious issue, as exemplified by the ongoing debates surrounding the licensing terms of models like Llama (Charlie Hull, 2023; Sriram Parthasarathy, 2024). The field is grappling with balancing open collaboration with intellectual property concerns and responsible AI development. Furthermore, the security of model hubs themselves has come under scrutiny. Services like model conversion tools hosted on platforms such as Hugging Face have been shown to be vulnerable to manipulation (Eoin Wickens, Kasimir Schulz, 2024), potentially allowing the introduction of malicious code into LLMs. The popularity of collaborative model merging techniques, while fostering innovation, also presents opportunities for introducing vulnerabilities (Hammoud et al., 2024).
@@
Challenge II [S&P + SE Perspective]: Risk Propagation in LLM Supply Chains. A significant challenge in the LLM supply chain lies in quantifying the propagation of risks across different models and components. The complex nature of LLMs, with their intricate architectures and numerous dependencies, makes it difficult to measure how vulnerabilities such as privacy breaches, biases, hallucinations, or potential backdoors propagate from one model to another. This challenge is exacerbated by the lack of standardized metrics and methodologies for assessing risk transmission. Model reuse techniques (Jiang et al., 2023), including fine-tuning, transfer learning, distillation, and model merging, add further complexity to this issue. When models are reused, there is a risk that vulnerabilities inherent in the original model can be unintentionally transferred to new iterations or applications. For instance, if a base model contains biases, these biases may be amplified or altered during the distillation process, leading to unforeseen consequences in downstream applications. Quantifying how such biases propagate through different stages of model reuse remains an open problem. Similarly, measuring the extent to which privacy vulnerabilities in training data persist through model iterations and affect derived models is not straightforward. The challenge is not only to identify these vulnerabilities but also to understand how they evolve as models are reused.
  • Opportunity: Developing Model Bill of Materials (MBOM). There lies a significant opportunity to enhance the security and integrity of LLMs through the development of standardized practices for generating and maintaining a MBOM for pre-trained models (David Mor-Ofek, 2024), mirroring the concept of SBOM. Such standardization would improve supply chain transparency, enabling stakeholders to more effectively identify, assess, and mitigate risks. Moreover, fostering collaboration among researchers, industry practitioners, and regulatory bodies can lead to the establishment of robust best practices and guidelines for the responsible release and sharing of models. This collaborative approach would not only enhance the trustworthiness and accountability of LLMs across the supply chain but also ensure that risk mitigation strategies are holistic, timely, and aligned with evolving ethical and security standards, ultimately leading to a safer and more reliable LLM ecosystem.
@@
Challenge IV [S&P Perspective]: Pre-trained Model Code Poisoning/Backdoor. The proliferation of model hubs as centralized platforms for sharing and distributing pre-trained models introduces significant security challenges. These hubs, while fostering collaboration and accessibility, also become potential points of failure and attractive targets for malicious actors. Key challenges include ensuring the integrity and authenticity of uploaded models, preventing the distribution of models with embedded malware or backdoors (Zhao et al., 2024b), and protecting against unauthorized access or data breaches. The dynamic nature of these platforms, with frequent updates and contributions from various sources, further complicates maintaining consistent security standards. Moreover, the potential for supply chain attacks, where compromised models could affect numerous downstream applications and users, presents a systemic risk to the LLM ecosystem.
  • Opportunity: Advanced Security Solutions for Model Hub Integrity. The security challenges posed by model hubs open up significant opportunities for developing robust security frameworks and technologies. There is potential to create advanced model verification systems that can automatically scan for malware, backdoors, or other vulnerabilities in uploaded models (Zhao et al., 2024b). Opportunities exist to implement measures to reduce model poisoning attack vectors. This could include developing sophisticated typo-squatting detection mechanisms to prevent malicious actors from exploiting similar or deregistered model names, organizations, or namespaces, addressing both typo-squatting and AIJacking vulnerabilities (Adrian Wood, 2023; Nadav Noy, 2023). Enhanced token management systems could mitigate the risk of API token leaks from prominent organizations (Bar Lanyado, 2023), thereby preventing unauthorized access and potential supply chain attacks. These advancements would not only enhance the security of individual model hubs but also contribute to building a more resilient and trustworthy LLM development ecosystem, reducing the risk of widespread supply chain attacks and unauthorized model access or manipulation (Eoin Wickens, Kasimir Schulz, 2024).